smith and vein had equal money. after smith spent 18 rupees and vein spent 25, smith had twice as much as vein. how much money did they had
Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

smith and vein had equal money. after smith spent 18 rupees and vein spent 25, smith had twice as much as vein. how much money did they had
-----
Equations:
s = v
-----
s-18 = 2(v-25)
-----
Substitute for "s" and solve for "v"::
v-18 = 2v - 50
v = 32 rupees
s = 32 rupees
---------------
how much money did they have?
Each had 32 rupees
----------
Cheers,
Stan H.
-------------

Answer by Boreal(15235)

x is what they started with
x-18=2(x-25)
x-18=2x-50
x=32 rupees
after spending 18, there are 14
after spending 25, there are 7